OpenStudy (anonymous):

can you explain how to solve tan^-1(sqrt3/3) in radians? i need help!!

OpenStudy (aum):

Let \(\tan^{-1}(\sqrt{3}/3) = x\). Find x in radians. \(\tan^{-1}(\sqrt{3}/3) = x ~~~~ \implies \tan(x) = \sqrt{3}/3\). Look up unit circle and determine for what angle in radians \(\tan(x) = \sqrt{3}/3\). That angle will be your x value.